# More variable circadian rhythms in epilepsy: a retrospective cross-sectional study using long-term heart rate recordings from wearable sensors

### v2

@@ 5 Code and Data Availability
6 AI statement
During the preparation of this work the authors used ChatGPT in order to improve the flow of the text and revise for grammar and typos. After using this tool/service, the authors reviewed and edited the content again and take full responsibility for the content of the publication.
7 Funding sources
Data collection for the study was supported by the Australian Government National Health and Medical Research Council (Grant #1178220). P.N.T. and Y.W. are both supported by UKRI Future Leaders Fellowships (MR/T04294X/1, MR/V026569/1).
@@ 8 Conflict of interest
The authors declare no conflict of interest.
@@
The simplest statistical modelling approach involves the fitting of a sinusoid wave. The rhythm produced is fixed in period, acrophase and amplitude. However, it can often be empirically observed (especially in uncontrolled conditions) that these properties vary over time from one circadian cycle to the next (9), potentially as the result of challenges to the circadian system (e.g inconsistent night shift working), variation in lifestyle and behaviour (e.g occasional late nights), or perhaps even disease (e.g the change between mood states in bipolar disorder, or seizure occurrence in epilepsy).
Bandpass-filtering of the signal around 24 hours derives a circadian rhythm where amplitude is able to vary across cycles, but the other parameters remain somewhat fixed. More sophisticated methods exist, such as wavelet-based analysis (the continuous and discrete wavelet transform), signal decomposition techniques (such as empirical mode decomposition) and others such as singular spectrum analysis (SSA) (Figure 2E(i)). An overview these methods can be found in (41). These methods are more flexible and produce a rhythm that fits the data better, capturing variation in all three properties across cycles. (12) evaluated these, as well as other, methods to extract a circadian rhythm from various physiological parameters, including heart rate. They compared the goodness-of-fit of the rhythm extracted by each method with the original data in each case. While certain methods performed better for some measures, SSA was the best all-rounder across measures.
@@
Missing data gaps in the recording were most likely caused by device charging or removal of the device by the participant for whatever other reason. As the heart rate recordings were computed by a proprietary Fitbit algorithm from raw PPG, it is possible that periods of noisy data (for example, due to poor device contact with skin) were excluded also. SSA, and other modelling methods, cannot handle missing data in the input recording, so interpolation (e.g linear) is required. However, there are occasionally very long (days-months) gaps of missing data for some participants, so we opted to avoid running the algorithm over very long interpolations of data, as this wasted computational resources and interfered with results. As such, the raw heart rate data was split into ‘runs’ between missing gaps larger than 8 hours. Within each run, any remaining gaps (which must be below 8 hours) were linearly interpolated.
@@
For each seven-day segment, the standard deviation of each property (Figure 2G) was calculated, reflecting the variability of that property during the corresponding week. Therefore, for each participant, a distribution of standard deviation values is produced for each property, reflecting differences in variability across weeks. To summarise these distributions for each participant, three summary values - circadian period variability, acrophase variability and amplitude variability - were calculated by taking the mean of the standard deviation values of each property across the segments distribution. Three additional summary values: period average, acrophase average and amplitude average, were calculated in a similar manner, using the mean of each property across all segments.
Calculating the mean and standard deviation separately for each segment ensures that they are always calculated on samples of data of the same length (7 cycles), reducing the problem of varying recording duration between participants. Taking the mean across all segments allows for comaprison of circadian average and variability between participants.
@@
Recording Duration mean (days) 221.7 443.9 1685 0.037
Recording Duration sd (days) 257.0 530.9
@@
Average Number of Seizures 42.7
@@
Period Mean -0.275 0.627
Period Std 2.597 0.009
Acrophase Mean 0.753 0.426
Acrophase Std 2.316 0.021
@@
Figure S5: Comparison of the distribution of the intra-individual average and variability of circadian properties between epilepsy types. Controls, mixed-type and unknown-type PWE are shown for reference only and are not included in the statistics shown. The two-sided Wilcoxon rank-sum test was used for comparison of the intra-individual average and variability between generalised and focal epilepsy. Each subpanel has been zoomed-in to omit outliers; outliers were not excluded prior to statistical testing.

### v4

@@ 6 Data Availability
7 Code availability
8 Acknowledgements
Data collection for the study was supported by the Australian Government National Health and Medical Research Council (Grant #1178220). P.N.T. and Y.W. are both supported by UK Research & Innovation (UKRI) Future Leaders Fellowships (MR/T04294X/1, MR/V026569/1).
9 Author contributions
BCS contributed to conceptualization, methodology, software, formal analysis, writing and visualisation; CT contributed to methodology, software, writing, visualisation, and supervision; RES contributed to resources, data curation, validation, and writing; GMB and NE contributed to validation and supervision; SJG contributed to validation and methodology; PNT contributed to writing, visualisation, supervision; KL contributed to methodology, validation, writing and supervision; PJK contributed to resources, data curation, validation, and writing; YW contributed to conceptualization, methodology, software, formal analysis, writing, visualisation, supervision, project administration, and funding acquisition. We also thank members of the Computational Neurology, Neuroscience & Psychiatry Lab (www.cnnp-lab.com) for discussions on the methodology, analysis and manuscript.
@@ 10 Conflict of interest
None of the authors have any conflict of interest to disclose. We confirm that we have read the Journal’s position on issues involved in ethical publication and affirm that this report is consistent with those guidelines.
@@
The simplest statistical modelling approach involves the fitting of a sinusoid wave. The rhythm produced is fixed in period, acrophase and amplitude. However, it can often be empirically observed (especially in uncontrolled conditions) that these properties vary over time from one circadian cycle to the next (9), potentially as the result of challenges to the circadian system (e.g inconsistent night shift working), variation in lifestyle and behaviour (e.g occasional late nights), or perhaps even health conditions (e.g the change between mood states in bipolar disorder, or seizure occurrence in epilepsy).
Bandpass-filtering of the signal around 24 hours derives a circadian rhythm where amplitude is able to vary across cycles, but the other parameters remain somewhat fixed. More sophisticated methods exist, such as wavelet-based analysis (the continuous and discrete wavelet transform), signal decomposition techniques (such as empirical mode decomposition) and others such as singular spectrum analysis (SSA) (Figure 2E(i)). An overview these methods can be found in (47). These methods are more flexible and produce a rhythm that fits the data better, capturing variation in all three properties across cycles.
(12) evaluated these as well as other methods for extraction of a circadian rhythm from a set of wearable measures (activity, heart rate, blood pressure, skin temperature, core temperature) recorded over one week from one individual. For each method and measure, they extracted a circadian rhythm and computed the goodness-of-fit ($R^{2}$) of the rhythm extracted compared to the original timeseries. While certain methods performed better on some measures, SSA was the best all-rounder across measures. We perform our own comparison using the same approach in Supplementary Figure S1, though with a smaller subset of methods and only using heart rate, but crucially we test over a much larger cohort (31 controls and 143 PWE). SSA performs best across methods (average $R^{2}\approx$ 0.4). Interestingly, there is a persistently lower $R^{2}$ across measures for PWE, potentially supportive of a more variable rhythm for PWE that is picked up poorly by less flexible methods (BANDPASS, COSINE), though could also be related to the increased sample size and recording duration of PWE - notably this difference is minimised for SSA. This finding further informs us that, on average, 40% of the variability in heart rate timeseries was accounted for by the circadian rhythm in this study, the remaining 60% associated with other factors, potentially exercise or acute stress.
While (12) found that certain methods (EEMD, CEEMDAN) performed better in heart rate, the sample size of 1 individual should be noted, as our result shows that, between participants, there is considerable variation and overlap in the $R^{2}$ between methods. For this reason, as well as SSA’s reliable performance in both (12) study and our own analysis, and our previous experience with these methods, we selected SSA for use in this study. Future work should compare a larger set of methods more thoroughly across a larger and more varied population.
Figure S1: The distribution of goodness-of-fit ($R^{2}$ (derived from Pearson’s r) averaged across runs for each participant) of the circadian rhythm extracted by each method compared to the participant’s original heart rate timeseries, split by PWE and controls. For methods requiring a minimum/maximum period range for the circadian rhythm the be detected within (CWT, BANDPASS), 20-30 hours were used respectively. This is the same threshold used to identify the circadian rhythm in the set of components extracted by SSA.
@@
Missing data gaps in the recording were most likely caused by device charging or removal of the device by the participant for whatever other reason. As the heart rate recordings were computed by a proprietary Fitbit algorithm from raw PPG, it is possible that periods of noisy data (for example, due to poor device contact with skin) were excluded automatically also. SSA cannot handle missing data in the input recording, so interpolation (e.g linear) is required. However, there are occasionally very long (days-months) gaps of missing data for some participants, so we opted to avoid running the algorithm over very long interpolations of data, as this wasted computational resources and interfered with results. As such, the raw heart rate data was split into ‘runs’ between missing gaps larger than 8 hours. Within each run, any remaining gaps (which as such must be below 8 hours) were linearly interpolated.
@@
For each seven-day segment, the standard deviation of each property (Figure 2G) was calculated, reflecting the variability of that property during the corresponding week. Therefore, for each participant, a distribution of standard deviation values is produced for each property, reflecting differences in variability across weeks. To summarise these distributions for each participant, three summary values - intra-individual circadian period variability, acrophase variability and amplitude variability - were calculated by taking the mean of the standard deviation values of each property across the segments distribution. Three additional summary values: intra-individual period average, acrophase average and amplitude average, were calculated in a similar manner, using the mean of each property across all segments.
Calculating the mean and standard deviation separately for each segment ensures that they are always calculated on samples of data of the same length (7 cycles), reducing the problem of varying recording duration between participants. Taking the mean across all segments allows for comparison of circadian average and variability between participants.
@@
Recording Duration mean (days) 221.5 443.9 1686 0.037
Recording Duration sd (days) 256.4 531.1
Recording Duration median (days) 125.1 215.8
Recording Duration IQR (days) 260.5 580.9
Recording Duration (min, max) (days) (7.0, 946.0) (6.7, 2715.6)
@@
Average Number of Seizures 43.4
@@
Period Mean -0.331 0.612
Period Std 2.569 0.01
Acrophase Mean 0.739 0.435
Acrophase Std 2.231 0.026
@@
Figure S6: Comparison of the distribution of the intra-individual average and variability of circadian properties between epilepsy types. Mixed type and unknown (not provided) type PWE are shown for reference only and are not included in the statistics shown. The two-sided Wilcoxon rank-sum test was used for comparison of the intra-individual average and variability between generalised and focal epilepsy, and each sub-type against controls. Each subpanel has been zoomed-in to omit outliers; outliers were not excluded prior to statistical testing.
S6 Impact of weekends and time of year on circadian properties and their intra-individual average and variability
To investigate whether variability in the circadian rhythm is associated with weekday vs. weekend effects or seasonal variations, circadian cycles (Figure 2G) were associated with a date (at acrophase), and 7-day segments were associated with their ‘time of year’ (see below).
S6.1 Weekday vs weekend effects
‘Social jetlag’ refers to the phenomenon where sleep patterns differ on weekdays compared to weekends due to the constraints of work, education or other commitments (8). More pronounced social jetlag has been reported for PWE compared to controls (37). Sleep patterns, while distinct from the circadian rhythm, are linked. As such, we test here whether circadian variability is similarly associated with weekday/weekend differences, and whether this varies between PWE and controls.
To compare intra-individual circadian average and variability between weekends and weekdays, each segment was split into ‘mid-week’ (Tue, Wed, Thu) and ‘circa-weekend’ (Fri, Sat, Sun) sub-segments, and intra-individual circadian average and variability values were calculated over mid-week and circa-weekend segments separately. The equal 3-day segment size was selected as a fully weekday/weekend (Mon-Fri vs Sat-Sun) comparison introduces a sample size bias (5 vs 2) that interferes with mean and standard deviation interpretation. Furthermore, we hypothesize this may account for behavioural shifts - individuals who experience social jetlag are most likely adjusted to their weekday schedule by Tuesday (following a disruptive Monday), and there is maybe a shift in behaviour on Friday evening compared to the evenings on days prior. Firstly, intra-individual circadian average and variability is compared between weekends and weekdays within PWE and control groups separately in Figure S7 to determine if a ‘weekend effect’ exists for both groups or is unique to either. Following this, Figure S8 compares intra-individual circadian average and variability between PWE and controls on weekdays and weekends separately, to determine if previously observed differences in these properties between PWE controls are consistent across the week, or are being driven by a weekend effect.
Figure S7: Intra-individual average and variability of circadian properties calculated over mid-week (Tue, Wed, Thu, lighter colours) and circa-weekend (Fri, Sat, Sun, darker colours) segments separately for both PWE and controls. The two-sided Wilcoxon signed-rank paired test was performed for comparison of the intra-individual average and variability calculated using mid-week vs circa-weekend segments for PWE and controls independently (i.e, is there a weekend effect within each group?). Only participants with at least 3 mid-week and at least 3 circa-weekend segments were kept. Highlighted p-values fall below the Bonferroni-corrected alpha threshold of 0.05/12 = 0.0042.
Figure S8: Intra-individual average and variability of circadian properties calculated over mid-week (Tue, Wed, Thu, lighter colours) and circa-weekend (Fri, Sat, Sun, darker colours) segments separately for both PWE and controls. The two-sided Wilcoxon rank-sum test was performed for comparison of the intra-individual average and variability calculated using mid-week vs circa-weekend segments separately, between controls and PWE (i.e, is there a weekend effect between groups?). Only participants with at least 3 mid-week and at least 3 circa-weekend segments were kept. Highlighted p-values fall below the Bonferroni-corrected alpha threshold of 0.05/12 = 0.0042.
To further explore the association between intra-individual average and variability of circadian properties in epilepsy and the weekend effect and determine the relative importance of these factors, a mixed-effects linear regression (MELR) model was used:
$\begin{split}property\_measure\sim&total\_duration\_d\\ &+condition\\ &+weekend\\ &+(condition:weekend)\\ &+(1|subject)\end{split}$ (1)
Where $property\_measure$ refers to Period Mean, Period Std, Acrophase Mean, etc, $total\_duration\_d$ refers to the total duration in days of all mid-week or circa-weekend segments for an individual, $condition$ refers to whether the individual is a control or PWE, $weekend$ refers to whether the $property\_measure$ is calculated over the mid-week (False) or circa-weekend (True) segments, and $(condition:weekend)$ refers to the interaction between the $condition$ and $weekend$ terms. A random intercept was applied for each participant to account for systematic differences between participants that may arise due to chronotype or routines (work, exercise, etc). The outputs of this model are shown in Table 2 with p-values $<0.05$ highlighted.
Additionally, to determine the relative importance of modelling the weekend effect, we performed likelihood ratio tests between the ‘full’ model (shown in Equation 1) and a ‘reduced’ model, that incorporates the weekend effect but not its interaction with condition (Equation 2) and a ‘null’ model that does not consider the weekend effect at all (Equation 3). p-values from these tests are reported in Table 3, with p-values $<$ 0.05 highlighted.
$\begin{split}property\_measure\sim&total\_duration\_d\\ &+condition\\ &+weekend\\ &+(1|subject)\end{split}$ (2)
$\begin{split}property\_measure\sim&total\_duration\_d\\ &+condition\\ &+(1|subject)\end{split}$ (3)
In Figure S7, most intra-individual average and variability measures differ between weekdays and weekends in PWE and controls. We avoid direct comparison of effect sizes or p-values here, as there are substantially different sample sizes underlying the PWE vs. controls cohorts. Instead we report them in Figure S7 “as is”.
In Figure S8, for each variability measure, there is a substantial difference in intra-individual variability of period and acrophase (Cohen’s d $\geq 0.6$), regardless of weekend or weekday. This confirms that the weekend effects observed in Fig S7 are not driving the results observed in Fig 3.
The MELR models outputs are reported in Table 2. Using the models to evaluate the interaction between the weekend and condition effects, there is weak evidence that intra-individual averages in Period and Amplitude are reduced on weekends only for PWE, although the associated p-values would not survive a correction for multiple comparison. No evidence of an interaction effect was seen in any of the intra-individual variability measures.
In Table 3, only Period and Amplitude Means have p-values below 0.05 in the reduced column, suggesting modelling the interaction between weekend and condition only can improve description of the data for these variables, which aligns with their entries in Table 2. However incorporation of at least the weekend effect alone is clearly necessary, as all p-values are <<much-less-than$<$ < 0.05 when comparing the null model (no incorporation of weekend effects) to the full.
Intra-individual circadian average and variability values appear to differ between weekdays and weekends for PWE and controls; but importantly, these differences do not drive the increased circadian variability we report in the main text. The weekday vs. weekend effect appear independently for both controls and PWE.
Period Mean Coef. Std.Err. z $P>|z|$ [0.025 0.975]
Intercept 24.005 0.042 570.531 0.000 23.922 24.087
C(Condition)[T.PWE] 0.073 0.045 1.607 0.108 -0.016 0.162
C(Weekend)[T.True] 0.013 0.055 0.243 0.808 -0.095 0.122
C(Condition)[T.PWE]:C(Weekend)[T.True] -0.122 0.060 -2.027 0.043 -0.240 -0.004
total_duration_d -0.000 0.000 -0.019 0.985 -0.000 0.000
Subject Var 0.004
Period Std Coef. Std.Err. z $P>|z|$ [0.025 0.975]
Intercept 0.860 0.072 11.994 0.000 0.719 1.000
C(Condition)[T.PWE] 0.240 0.077 3.119 0.002 0.089 0.391
C(Weekend)[T.True] 0.105 0.062 1.712 0.087 -0.015 0.226
C(Condition)[T.PWE]:C(Weekend)[T.True] 0.019 0.067 0.290 0.772 -0.112 0.151
total_duration_d -0.000 $0$1.364 0.173 -0.000 0.000
Subject Var 0.072 0.073
Acrophase Mean Coef. Std.Err. z $P>|z|$ [0.025 0.975]
Intercept 15.312 0.470 32.555 0.000 14.390 16.233
C(Condition)[T.PWE] 0.137 0.504 0.273 0.785 -0.850 1.125
C(Weekend)[T.True] 0.265 0.113 2.349 0.019 0.044 0.487
C(Condition)[T.PWE]:C(Weekend)[T.True] 0.076 0.123 0.621 0.535 -0.165 0.318
total_duration_d -0.001 0.001 -0.635 0.525 -0.002 0.001
Subject Var 4.884 2.063
Acrophase Std Coef. Std.Err. z $P>|z|$ [0.025 0.975]
Intercept 0.550 0.052 10.651 0.000 0.449 0.651
C(Condition)[T.PWE] 0.178 0.055 3.204 0.001 0.069 0.286
C(Weekend)[T.True] 0.072 0.046 1.555 0.120 -0.019 0.162
C(Condition)[T.PWE]:C(Weekend)[T.True] -0.025 0.050 -0.494 0.621 -0.123 0.074
total_duration_d -0.000 $0$1.645 0.100 -0.000 0.000
Subject Var 0.036 0.049
Amplitude Mean Coef. Std.Err. z $P>|z|$ [0.025 0.975]
Intercept 11.153 0.683 16.342 0.000 9.816 12.491
C(Condition)[T.PWE] -0.549 0.731 -0.751 0.453 -1.981 0.884
C(Weekend)[T.True] 0.411 0.236 1.742 0.082 -0.051 0.873
C(Condition)[T.PWE]:C(Weekend)[T.True] -0.543 0.257 -2.114 0.035 -1.047 -0.040
total_duration_d -0.001 0.001 -0.460 0.645 -0.003 0.002
Subject Var 9.928 2.045
Amplitude Std Coef. Std.Err. z $P>|z|$ [0.025 0.975]
Intercept 1.459 0.099 14.730 0.000 1.265 1.653
C(Condition)[T.PWE] 0.079 0.106 0.748 0.455 -0.129 0.288
C(Weekend)[T.True] 0.222 0.080 2.790 0.005 0.066 0.379
C(Condition)[T.PWE]:C(Weekend)[T.True] -0.085 0.087 -0.982 0.326 -0.256 0.085
total_duration_d -0.000 0.000 -0.814 0.415 -0.001 0.000
Subject Var 0.148 0.110
Table 2: Summary of the mixed effects models produced by Equation 1.
reduced null
Period Mean 0.032 0.0
Period Std 0.772 0.0
Acrophase Mean 0.535 0.0
Acrophase Std 0.621 0.021
Amplitude Mean 0.036 0.098
Amplitude Std 0.327 0.0
Table 3: p-values for likelihood ratio tests comparing reduced (Equation 2) and null (Equation 3) models to the full (Equation 1) model for each property measure. p $<$ 0.05 reject the null hypothesis and suggests the full model fits the data significantly better and should be used.
S6.2 Time of year effects
Seizures have been shown to exhibit seasonal variation, possibly clustering in winter (48), and seizure occurrences have not only been found to be phase-locked to circadian rhythms of heart rate for some PWE, but longer-term (weekly and monthly) rhythms of heart rate also (14). Here, we test whether intra-individual variability is associated with long-term seasonal trends, and whether this varies between PWE and controls.
To explore changes in intra-individual circadian average and variability over the year, 7-day segments (Figure 2G) were associated with the ‘time of the year’ ($toy$) at which they occurred. This is defined as $toy=\frac{day\_of\_year}{days\_in\_year}$ where $day\_of\_year$ is an integer 1-365 (e.g. 1=1st January, 365=31st December) corresponding to the 4th (middle) circadian cycle of each 7-day segment, and days_in_year is 366 on leap years or 365 on regular years. Rather than producing a summary statistic for each participant by calculating the mean of circadian average and variability values across all segments as before, we can investigate whether these individual segment values vary depending on the time of year of the segment. This allows for a higher resolution in time along the year, and this is visualised for an example participant in Figure S9.
Figure S9: A visualisation of the data used in the MELR for one example participant. Each point corresponds to the circadian average or variability value for one 7-day segment, placed corresponding to its position within the year (data may span multiple years). Acrophase means have been corrected for local timezone and daylight saving time shifts. Coloured regions correspond to the seasons (Australian): Spring (Sept-Nov, pink), Summer (Dec-Feb, green), Autumn (Mar-May, orange), and Winter (Jun-Aug, blue).
To model the association between intra-individual average and variability of circadian properties in epilepsy and annual variations and determine the relative importance of these factors, a mixed-effects linear regression model (MELR) was used:
$\begin{split}property\_measure\sim&condition\\ &+sin(toy\_radians)\\ &+cos(toy\_radians)\\ &+(condition:(sin(toy\_radians)+cos(toy\_radians)))\\ &+(1|subject)\end{split}$ (4)
Where $property\_measure$ refers to Period Mean, Period Std, Acrophase Mean, etc, of a given segment, and $condition$ refers to whether the individual is a control or PWE. As the time of year is a circular variable, it cannot be used directly in a linear regression model. To account for this, cos and sin of the month variable (after conversion to radians ($toy\_radians=toy\cdot 2\pi$)) were included. A significant association with either one or both of these terms can be interpreted as a seasonal effect being present. The interaction between these terms and condition was added via $condition:(sin(toy\_radians)+cos(toy\_radians))$. A random intercept was applied for each participant to account for systematic differences between participants that may arise due to chronotype or routines (work, exercise, etc); but here, also to account for different number of 7-day segments between subjects. The outputs of this model are shown in Table 4 with p-values $<$ 0.05 highlighted.
Additionally, to determine the relative importance of modelling the time of year effect, we performed likelihood ratio tests between the ‘full’ model (shown in Equation 4) and a ‘reduced’ model, that incorporates the time of year effect but not its interaction with condition (Equation 5) and a ‘null’ model that does not consider the time of year effect at all (Equation 6). p-values from these tests are reported in Table 5 with p-values $<$ 0.05 highlighted.
$\begin{split}property\_measure\sim&condition\\ &+sin(toy\_radians)\\ &+cos(toy\_radians)\\ &+(1|subject)\end{split}$ (5)
$\begin{split}property\_measure\sim&condition\\ &+(1|subject)\end{split}$ (6)
Period Mean Coef. Std.Err. z $P>|z|$ [0.025 0.975]
Intercept 24.038 0.024 989.150 0.000 23.990 24.085
C(Condition)[T.PWE] 0.016 0.026 0.611 0.541 -0.036 0.068
sin_toy_radians 0.027 0.018 1.518 0.129 -0.008 0.061
C(Condition)[T.PWE]:sin_toy_radians -0.038 0.019 -2.022 0.043 -0.074 -0.001
cos_toy_radians -0.006 0.019 -0.341 0.733 -0.043 0.031
C(Condition)[T.PWE]:cos_toy_radians 0.014 0.020 0.715 0.475 -0.025 0.053
Subject Var 0.009 0.004
Period Std Coef. Std.Err. z $P>|z|$ [0.025 0.975]
Intercept 1.011 0.072 14.079 0.000 0.871 1.152
C(Condition)[T.PWE] 0.265 0.079 3.368 0.001 0.111 0.419
sin_toy_radians -0.036 0.024 -1.531 0.126 -0.083 0.010
C(Condition)[T.PWE]:sin_toy_radians 0.024 0.025 0.966 0.334 -0.025 0.073
cos_toy_radians -0.005 0.025 -0.210 0.833 -0.055 0.045
C(Condition)[T.PWE]:cos_toy_radians -0.014 0.027 -0.537 0.591 -0.067 0.038
Subject Var 0.127 0.030
Acrophase Mean Coef. Std.Err. z $P>|z|$ [0.025 0.975]
Intercept 15.394 0.405 37.978 0.000 14.599 16.188
C(Condition)[T.PWE] 0.031 0.446 0.069 0.945 -0.843 0.905
sin_toy_radians 0.076 0.066 1.149 0.250 -0.053 0.205
C(Condition)[T.PWE]:sin_toy_radians -0.104 0.069 -1.504 0.132 -0.240 0.032
cos_toy_radians 0.084 0.070 1.194 0.232 -0.054 0.222
C(Condition)[T.PWE]:cos_toy_radians 0.231 0.074 3.141 0.002 0.087 0.376
Subject Var 4.750 0.391
Acrophase Std Coef. Std.Err. z $P>|z|$ [0.025 0.975]
Intercept 0.867 0.082 10.616 0.000 0.707 1.027
C(Condition)[T.PWE] 0.252 0.090 2.815 0.005 0.077 0.427
sin_toy_radians -0.017 0.026 -0.662 0.508 -0.069 0.034
C(Condition)[T.PWE]:sin_toy_radians -0.012 0.028 -0.436 0.663 -0.066 0.042
cos_toy_radians -0.009 0.028 -0.313 0.754 -0.064 0.046
C(Condition)[T.PWE]:cos_toy_radians -0.026 0.029 -0.873 0.383 -0.084 0.032
Subject Var 0.166 0.035
Amplitude Mean Coef. Std.Err. z $P>|z|$ [0.025 0.975]
Intercept 10.951 0.592 18.511 0.000 9.791 12.110
C(Condition)[T.PWE] -0.692 0.651 -1.062 0.288 -1.968 0.585
sin_toy_radians 0.289 0.096 3.010 0.003 0.101 0.477
C(Condition)[T.PWE]:sin_toy_radians -0.024 0.101 -0.239 0.811 -0.222 0.174
cos_toy_radians 0.678 0.103 6.603 0.000 0.476 0.879
C(Condition)[T.PWE]:cos_toy_radians -0.335 0.107 -3.122 0.002 -0.546 -0.125
Subject Var 10.119 0.541
Amplitude Std Coef. Std.Err. z $P>|z|$ [0.025 0.975]
Intercept 2.013 0.114 17.689 0.000 1.790 2.236
C(Condition)[T.PWE] -0.027 0.125 -0.216 0.829 -0.271 0.217
sin_toy_radians 0.010 0.039 0.268 0.788 -0.066 0.086
C(Condition)[T.PWE]:sin_toy_radians 0.011 0.041 0.266 0.790 -0.069 0.091
cos_toy_radians 0.100 0.042 2.416 0.016 0.019 0.182
C(Condition)[T.PWE]:cos_toy_radians -0.049 0.043 -1.136 0.256 -0.135 0.036
Subject Var 0.316 0.046
Table 4: Summary of the mixed effects models produced by Equation 4.
reduced null
Period Mean 0.091 0.106
Period Std 0.526 0.028
Acrophase Mean 0.002 0.0
Acrophase Std 0.634 0.0
Amplitude Mean 0.008 0.0
Amplitude Std 0.497 0.0
Table 5: p-values for likelihood ratio tests comparing reduced (Equation 5) and null (Equation 6) models to the full (Equation 4) model for each property measure. p $<$ 0.05 reject the null hypothesis and suggests the full model fits the data significantly better and should be used.
For the individual in Figure S9, period and acrophase averages are consistent over the course of the year, but other properties display seasonal variation.
Looking at the MELR output (Table 4), some seasonal effects were observed for intra-individual average and variability in PWE and controls, most pronounced in intra-individual average amplitude. However, we found no evidence that the reported increase in period and acrophase variability in PWE were driven by seasonal variations.
In Table 5, all rows bar average period have a p-value below 0.05 when compared to the null model, indicating that consideration of seasonal effects helps explain some variance in our circadian properties. Only average acrophase and amplitude had p-values below 0.05 in the reduced column, implying there may be some interaction between epilepsy and annual variation in intra-individual average acrophase and amplitude.
Intra-individual circadian average and variability values appear to vary subtly over the year in both PWE and controls. However, the relevance of this to epilepsy is not as clear: LR testing identified that average acrophase and amplitude are possibly altered in epilepsy, which could imply that PWE experience more of a seasonal shift in these properties than controls, but this needs to be explored more in future studies with substantially more controls. Most important to this work, we found no evidence of seasonal effects in intra-individual variability of period and acrophase. Thus our main results are unlikely to be driven by seasonal effects.